Commercial aircraft used for flying in instrument conditions are required to have two independent radios instead of one. AssuWhy does the usual rounding rule of three significant digits not work here? A. Rounding the result to three significant digitIs this probability high enough to ensure flight safety? A. No, the probability is not high enough to ensure flight safety beAssume that the significance level is a = 0.05. Use the given information to find the P-value and the critical value(s). TheThe critical value(s) is/are z= (Round to two decimal places as needed. Use a comma to separate answers as needed.)
Answer #1

Answer)

Answer)

As there are fixed number of trials and probability of each and every trial is same and independent of each other

Here we need to use the binomial formula

P(r) = ncr*(p^r)*(1-p)^n-r

Ncr = n!/(r!*(n-r)!)

N! = N*n-1*n-2*n-3*n-4*n-5........till 1

For example 5! = 5*4*3*2*1

Special case is 0! = 1

P = probability of single trial = 1 - 0.0023 = 0.9977

N = number of trials = 2

R = desired success = at least 1.

P(at least 1) = 1 - P(0) = 0.99999471

C. Rounding the result to three significant digits yields 1.00, which is misleading because it suggests that it is certain that both radios will work correctly.

Part 2)

From z table, P(z>1.34) = 0.0901

P-value = 0.0901

From z table, P(z>1.645) = 0.05

So, critical value is 1.645